# Harnessing the Universal Geometry of Embeddings

### v3

@@ Appendix B Extended oracle-aided optimal transport baseline results
using four algorithms: Hungarian (linear sum assignment), Earth Mover’s Distance (EMD), Sinkhorn, and Gromov-Wasserstein. Note that this baseline computes matchings and transports between embeddings derived from the same underlying texts, strongly favoring optimal transport (OT) methods. Nevertheless, OT still struggles when embeddings originate from different model backbones.
@@
Appendix C Full out-of-distribution translation results
@@
$E_{1}$ $E_{2}$ $\cos(\cdot)$ T-1 Rank $\cos(\cdot)$ T-1 Rank $\cos(\cdot)$ T-1 Rank
gra. gtr 0.74 (0.0) 0.99 1.09 (0.1) -0.04 (0.0) 0.00 415.61 (8.2) 0.51 (0.0) 0.01 398.32 (8.2)
gte 0.85 (0.0) 0.95 1.26 (0.1) 0.00 (0.0) 0.00 406.73 (8.2) 0.74 (0.0) 0.01 398.14 (8.2)
ste. 0.77 (0.0) 0.96 1.11 (0.0) 0.00 (0.0) 0.00 417.27 (8.2) 0.56 (0.0) 0.00 399.57 (8.2)
e5 0.83 (0.0) 0.87 3.10 (0.7) 0.02 (0.0) 0.00 405.53 (8.1) 0.75 (0.0) 0.01 398.33 (8.2)
@@
Table 6: Out-of-distribution translations on TweetTopic (with baselines): vec2vec models trained on NQ and evaluated on the entire TweetTopic test set (800 tweets). The rank metric varies from 1 to 800, thus 400 corresponds to a random ordering. Standard errors are shown in parentheses. Symbols denote the lowest-rank solver: Earth Mover’s Distance, Sinkhorn and Gromov-Wasserstein
@@
$E_{1}$ $E_{2}$ $\cos(\cdot)$ T-1 Rank $\cos(\cdot)$ T-1 Rank $\cos(\cdot)$ T-1 Rank
@@
Appendix D Ablations
D.1 Ablating components of our method
We ablate our method subtractively, measuring the key metrics after removing individual components of our algorithm (described in Section˜3). Table˜8 shows that each component appears to be critical to building good translations. While in each setting, vec2vec’s $\cos(\cdot)$ is higher than the naïve baselines, the Top-1 accuracies and ranks imply that ablated vec2vec translations are at best only slightly better than the baseline and do not preserve the geometry of the vector space.
Method $\cos(\cdot)$ Top-1 Rank
vec2vec 0.75 (0.0) 0.91 2.64 (0.1)
Naïve Baseline 0.04 (0.0) 0.00 4084.15 (9.2)
OT Baseline 0.70 (0.0) 0.00 4078.45 (9.2)
– VSP loss 0.58 (0.0) 0.00 4196.64 (9.2)
– CC loss 0.50 (0.0) 0.00 3941.36 (9.3)
– latent GAN 0.49 (0.0) 0.00 3897.09 (9.5)
– VSP and CC loss 0.47 (0.0) 0.00 3365.24 (9.3)
– hyperparam. tuning 0.50 (0.0) 0.00 4011.73 (9.3)
D.2 Amount of data needed to learn translation
NN $\cos(\cdot)$ Top-1 Rank
1000000 0.75 (0.0) 0.92 2.73 (0.2)
10000 0.57 (0.0) 0.01 1462.21 (20.)
50000 0.74 (0.0) 0.81 3.91 (0.6)
100000 0.74 (0.0) 0.85 4.52 (0.4)
500000 0.75 (0.0) 0.92 2.73 (0.2)
In our main experiments, we use 1M-point subsets of NQ to train our vec2vec models. Now, we train the gtr-gte vec2vec with 1M gtr embeddings but fewer gte embeddings. Table˜9 shows that translators still learn something (i.e. are better than random), and only 50K embeddings are sufficient to learn a translation that is almost as good. Even with 10K embeddings, our method still learns a non-trivial translation. Translation performance generally seems to improve with the amount of training data, however.
Appendix E Prompt for measuring information extraction
Figure˜5 utilizes an LLM judge to determine whether the approximate inversion (using [66]) of a translated email embedding leaks information about the original email. Figure˜7 shows our prompt to the GPT-4o judge:

### v4

@@ Appendix B Oracle-aided optimal transport baseline
using four algorithms: Hungarian (linear sum assignment), Earth Mover’s Distance (EMD), Sinkhorn, Gromov-Wasserstein. For the Gromov-Wasserstein algorithm, we try both the entropic and non-entropic variants with multiple hyperparameter configurations and select the best figure. Note that the optimal transport (OT) baseline computes matchings and transports between embeddings derived from the same underlying texts, strongly favoring OT methods. Nevertheless, OT still struggles when embeddings originate from different model backbones.
@@
Appendix C Translating to and from Qwen
vec2vec OT Baseline
gte qwen 0.50 (0.0) 0.92 2.28 (0.2) 0.38 (0.0) 0.00 425.28 (1.1)
qwen gte 0.84 (0.0) 0.88 2.49 (0.3) 0.85 (0.0) 0.00 425.07 (1.2)
Table 8: Translations between GTE and Qwen embeddings trained on NQ and evaluated on a 65536 text subset of NQ (chunked in batches of size 1024). Rank varies from 1 to 1024, thus 512 corresponds to a random ordering. Since the embedding dimensionalities are different, only the Gromov-Wasserstein OT baseline is run and the naive baseline does not apply. Bold denotes best value.
[…]
Appendix D Text-image retrieval on MS COCO
model R@16 $\uparrow$ $\cos(\cdot)$ $\uparrow$ Rank $\downarrow$
granite $\to$ clip 0.23 0.23 (0.0) 233.67 (3.0)
clip (baseline) 0.75 0.30 (0.0) 23.20 (0.8)
Table 9: Cross-model text-image retrieval on MS COCO: granite $\to$ clip vec2vec trained on NQ (unimodal) and evaluated on MS COCO’s validation set. The rank metric varies from 1 to 5000, thus 2500 corresponds to a random ordering. Queries (captions) embedded with either Granite or CLIP. Documents (images) embedded with CLIP. Each caption has a unique image. Standard errors are shown in parentheses.
Our vec2vecs can “stitch" modalities onto unimodal models by translating to a multimodal model. To test this, we evaluated cross-modal text-image retrieval on MS COCO’s validation set (5000 examples) [undefag], translating queries (captions) embedded with Granite to retrieve documents (images) embedded with CLIP using our unimodal granite $\to$ clip translator from section˜4.3. Each caption has a unique image. We report Recall@16, cosine similarities, and Rank, with CLIP (for both documents and queries) as our baseline.
As Table˜9 shows, translating Granite embeddings to CLIP enables non-negligible cross-model multimodal retrieval with a unimodal model for queries—despite zero multimodal training. Further evaluation of this paradigm with multimodal-specific training is a promising direction.
Appendix E Initialization robustness by model backbone
GAN training is notoriously unstable to weight initialization [undefaz]. To measure our method’s robustness, we trained fifteen e5 $\to$ gte (shared backbone) and e5 $\to$ gtr (cross-backbone) vec2vecs on the NQ dataset for a fixed 10 epochs.
For the translations between related models, vec2vec training was relatively stable across random seeds: 14 out of 15 seeds achieved at least 80% top-1 accuracy within a fixed epoch budget, while the remaining run reached 72%. In contrast, translation between unrelated models proved significantly less stable, with only 3 out of 15 runs achieving convergence (80% top-1 accuracy). We leave improving the seed stability of our training regime as future, valuable work.
Appendix F Full out-of-distribution translation results
@@
$E_{1}$ $E_{2}$ $\cos(\cdot)$ $\uparrow$ T-1 $\uparrow$ Rank $\downarrow$ $\cos(\cdot)$ $\uparrow$ T-1 $\uparrow$ Rank $\downarrow$ $\cos(\cdot)$ $\uparrow$ T-1 $\uparrow$ Rank $\downarrow$
gra. gtr 0.74 (0.0) 0.99 1.09 (0.1) -0.04 (0.0) 0.00 415.61 (8.2) 0.71 (0.0) 0.01 220.93 (7.1)
gte 0.85 (0.0) 0.95 1.26 (0.1) 0.00 (0.0) 0.00 406.73 (8.2) 0.87 (0.0) 0.01 201.48 (6.6)
ste. 0.77 (0.0) 0.96 1.11 (0.0) 0.00 (0.0) 0.00 417.27 (8.2) 0.74 (0.0) 0.00 239.36 (6.7)
e5 0.83 (0.0) 0.87 3.10 (0.7) 0.02 (0.0) 0.00 405.53 (8.1) 0.87 (0.0) 0.01 244.94 (7.4)
@@
Table 10: Out-of-distribution translations on TweetTopic (with baselines): vec2vec models trained on NQ and evaluated on the entire TweetTopic test set (800 tweets). The rank metric varies from 1 to 800, thus 400 corresponds to a random ordering. Standard errors are shown in parentheses. Symbols denote the lowest-rank solver: Earth Mover’s Distance and Gromov-Wasserstein
@@
$E_{1}$ $E_{2}$ $\cos(\cdot)$ $\uparrow$ T-1 $\uparrow$ Rank $\downarrow$ $\cos(\cdot)$ $\uparrow$ T-1 $\uparrow$ Rank $\downarrow$ $\cos(\cdot)$ $\uparrow$ T-1 $\uparrow$ Rank $\downarrow$
@@
Appendix G Zero-shot inversion on TweetTopic
Figure 7: Leakage of information via inversion. Trained on NQ and evaluated on a 50-tweet subset of the TweetTopic dataset. Cells denote judge accuracy.
We replicate Figure˜5 on the TweetTopic dataset in Figure˜7, extracting information given only translated embeddings of tweets. We achieve non-negligible leakage for all model pairs.
Appendix H Prompt for measuring information extraction
Figures˜5 and 7 utilize an LLM judge to determine whether the approximate inversion (using [undefaan]) of a translated email embedding leaks information about the original email. Figure˜8 shows our prompt to the GPT-4o judge: